40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 811     Entry time: Thu Aug 7 17:32:23 2008
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Afternoon PRM activities 
Rana, Jenne, Yoichi, Dmass

After Yoichi confirmed this morning that the wire was in both grooves, Rana attempted to lift the PRM a tiny bit, and twist it around (very gently of course) to see if we could make the wire slip back to its nominal position underneath the optic. On the first attempt, the wire ended up slipping the wrong way, causing slightly more tilt. On another attempt, the wire came out of the groove nearest the chamber door by about 0.5mm. We got the wire back in the groove by slightly lifting the optic, and pushing the wire back in. Then, on further attempts at making the wire slip back to its nominal position, the wire came out of the groove farthest from the chamber door. It is very difficult to get at that side of the PRM, because the table is crowded, and it is on the far side of the optical table from the chamber door. We decided to pull the PRM out of the chamber. Rana clamped the mirror into its cage using the earthquake stops and removed the OSEMS, and then we pulled the mirror out. We put it on a cart that was covered with foil and had a little foil house for the optic cage. We rolled the mirror+cage over to the flow bench at the end of the y-arm.

We saw that the wire is no longer even on the standoff (~3mm away from the groove) on the side that was farthest from the chamber door.

Since we have not confirmed that we have spare wire and spare magnets (and due to the time of day), we have decided to cover the cage with some foil, while it is sitting on the flow bench, and we'll fix the wire in the morning.
ELOG V3.1.3-